# Machine-Made Media: Monitoring the Mobilization of Machine-Generated Articles on Misinformation and Mainstream News Websites

### 17 Jan 2024 v4

@@ Abstract
As large language models (LLMs) like ChatGPT have gained traction, an increasing number of news websites have begun utilizing them to generate articles. However, not only can these language models produce factually inaccurate articles on reputable websites but disreputable news sites can utilize LLMs to mass produce misinformation. To begin to understand this phenomenon, we present one of the first large-scale studies of the prevalence of synthetic articles within online news media. To do this, we train a DeBERTa-based synthetic news detector and classify over 15.90 million articles from 3,074 misinformation and mainstream news websites. We find that between January 1, 2022, and May 1, 2023, the relative number of synthetic news articles increased by 55.4% on mainstream websites while increasing by 457% on misinformation sites. We find that this increase is largely driven by smaller less popular websites. Analyzing the impact of the release of ChatGPT using an interrupted-time-series, we show that while its release resulted in a marked increase in synthetic articles on small sites as well as misinformation news websites, there was not a corresponding increase on large mainstream news websites.
@@ 1 Introduction
In this work, we present a large-scale study of the relative increase in machine-generated/synthetic articles from 3,074 news websites (1,114 misinformation/unreliable websites and 1,960 mainstream/reliable news websites) between January 1, 2022, and May 1, 2023. To do this, we utilize data from 19 models, as well as adversarial data from article perturbation/re-writes and paraphrases, to train a DeBERTa-based model (He et al. 2021) to detect English-language synthetic news articles. We subsequently benchmark this classifier on eight test sets of machine-generated news articles, including two from real-world companies (Pu et al. 2022) and one from an independent collection of human-written real-world articles. Across these test datasets, our model, at a false positive rate (FPR) of 1%, achieves an average precision score of 0.992. With this trained model, we classify over 15.90M articles published between January 1, 2022, and May 1, 2023, from our set of 3,074 news websites. [^1] We release the weights of our model and the URLs used in this study at https://github.com/hanshanley/machine-made-media.
We find that among reliable/mainstream news websites, synthetic articles increased in prevalence by 55.4% (0.91% of news articles in January 2022 to 1.42% in May 2023) while among unreliable/misinformation websites, the prevalence increased by 457% (0.38% of news articles in January 2022 to 2.14% in May 2023). Examining the content of synthetic articles, we find that while mainstream/reliable news websites have largely utilized synthetic articles to report on financial and business-related news, misinformation/unreliable news websites have reported on a wide range of topics ranging from world affairs (e.g., the Russo-Ukrainian War) to human health (e.g., COVID-19). Examining the impact of ChatGPT on the prevalence of synthetic content, we further find that its release coincided with significant increases in machine-generated articles on misinformation websites and unpopular mainstream news websites.
@@ 3 Detecting Machine-Generated Articles
We finally note that, as found in prior work (Pu et al. 2022; Uchendu et al. 2021; Zellers et al. 2019), machine-generated news articles are often shorter in length than human-written articles. While training, to ensure that our models do not simply distinguish between longer human-written articles and those generated by generative transformers by their different lengths, we ensure that our machine-generated and human-written articles are of similar lengths (median training synthetic article length of 210 words and median training human article length of 224 words). Furthermore, as found by past work, predictions for texts, particularly short texts, tend to be unreliable (Kirchner et al. 2023; Zellers et al. 2019; Pu et al. 2022); conversely, as shown by Sadasivan et al. (Sadasivan et al. 2023), as the lengths of texts increase the variance between human and machine-generated texts increases. As such, for our training and our generated test data (GPT 3.5 dataset), we exclude texts shorter than 1,000 characters (140 words) (OpenAI 2022). We note as a result, we do not use every trained model’s articles from the Turing Benchmark; given that WMT-20/FAIR articles within this dataset are all shorter than 1000 characters, we do not include them within our training dataset. Altogether our training dataset thus includes data from 19 different models (18 from Turing Benchmark and GPT-3.5 davinici).
@@
Constructing Paraphrased Synthetic Articles. To paraphrase each of the machine-generated articles within our dataset, we use the approach outlined by Krishna et al. (Krishna et al. 2023). Specifically, as in their work, we utilize Dipper, a version of the T5 generative model fine-tuned on paragraph-level paraphrases, that outputs paraphrased versions of the inputted text. We use the default and recommended parameters [^5] We use a lexical diversity parameter of 60. For more details on the Dipper model see Krishna et al. (Krishna et al. 2023) as in Krishna et al. to paraphrase a portion of the text within our original training dataset as well as our GPT-3.5 test dataset (Krishna et al. 2023). We note that after paraphrasing our datasets, we again filter to ensure all articles utilized for training contain at least 1000 characters (Table 2). We annotate training and test datasets containing articles paraphrased with Dipper with the suffix Para. After paraphrasing we still consider these articles to be synthetic.
@@ 4 News Dataset and Classification Pipeline
Website List. Between January 1, 2022, and May 1, 2023, we gather all articles published from 3,074 news websites. [^10] We note that while this study focuses on the release of ChatGPT as a possible focal point, our data collection for this project actually began in January 2022. Our list of websites consists of domains labeled as “news” by Media Bias Fact Check [^11] https://mediabiasfactcheck.com/ and by prior work (Hanley, Kumar, and Durumeric 2023). Within our list of news sites, we differentiate between “unreliable news websites” and “reliable news websites.” Our list of unreliable news websites includes 1,114 domains labeled as “conspiracy/pseudoscience” by mediabiasfactcheck.com as well as those labeled as “unreliable news”, misinformation, or disinformation by prior work (Hanley, Kumar, and Durumeric 2023; Barret Golding 2022; Szpakowski 2020). Our set of “unreliable” or misinformation news websites includes websites like realjewnews.com, davidduke.com, thegatewaypundit.com, and breitbart.com. We note that despite being labeled unreliable every article from each of these websites is not necessarily misinformation.
Our set of “reliable” news websites consists of the news websites that were labeled as belonging to the “center”, “center-left”, or “center-right” by Media Bias Fact Check as well as websites labeled as “reliable” or “mainstream” by other works (Hanley, Kumar, and Durumeric 2023; Barret Golding 2022; Szpakowski 2020). This set of “reliable news websites” includes websites like washingtonpost.com, reuters.com, apnews.com, cnn.com, and foxnews.com. Altogether after removing duplicates and unavailable websites, we scraped 1,960 “reliable news” or mainstream websites.
@@
Given that many of our websites (e.g., cnn.com) have multilingual options, we use the Python langdetect library to filter out all non-English articles. To prepare data for classification, we remove boilerplate language using the Pythonjustext library and then remove URLs, emojis, and HTML tags. Further, to ensure the reliability of our classifications, we only classify news articles that are at least 1000 characters (approximately 140 words) long. Altogether, from our selection of 3,074 websites, we gathered 15.90M articles that were published between January 1, 2022, and May 1, 2023. Finally, we utilize our DeBERTa+Pert+Para model at a softmax classification threshold of 0.98 to classify each article as either human-written or machine-generated. Classifying all 15.90M articles took approximately 65.8 hours using an Nvidia RTX A6000 GPU.
@@
With the rise of LLMs, many companies have widely scraped and gathered data from websites to fuel their models (Schappert 2023). As a result, websites ranging from Twitter to Reddit have begun to set up restrictions to ensure the privacy of their users and to protect their content from being used in other private companies’ generative models. While we do not train a generative model that could artificially produce convincing and seemingly unique reproductions of the texts that we utilize, we note the concern that our work raises. Our work, however, only studies the texts of our set of 15.90 million articles and classifies them as machine-generated or human-written. We do not seek to generate summaries or artificial rewrites of this content. In terms of web crawling for this data, as noted elsewhere (Singrodia, Mitra, and Paul 2019; Hanley, Kumar, and Durumeric 2023; Smith et al. 2013), website crawling and scraping remain pivotal for understanding and documenting what occurs on the Internet. Without scraping, understanding trends and how the Internet could potentially affect real life becomes impossible. As decided in Van Burn v. United States, publically accessible information can be legally scraped as long as it is done ethically and does no harm to the site (Emily R. Lowe and Katrina Slack 2022). As such, we collect only publicly available data from our set of websites and follow the best practices for web crawling as in Acar et al. (2014). We limit the load that each news site experiences by checking for new articles daily at a maximum rate of one request every 10 seconds. The hosts that we scan from are identifiable through WHOIS, reverse DNS, and an HTTP landing page explaining how to reach us if they would like to be removed from the study. During our crawling period, we received no requests from websites to opt out.
@@ 5 The Rise of Machine-Generated Media
Large-Scale Trends in Machine-Generated Media. To begin, we plot the average percentage of synthetic news articles per website across our dataset between January 1, 2022, and May 1, 2023, in Figure 1. In aggregate, across all 3,074 sites, we see that 1.07% of all articles published in January 2022 (13,447 of 1,254,959) were synthetically generated. However, by May 2023, the fraction of synthetic articles nearly went up to 1.76% (26,006 of 1,469,612 articles), a 65.1% relative increase (nearly doubling in raw amount).
We observe that our set of reliable/mainstream websites typically had a greater percentage of synthetic articles at the beginning of 2022 compared with misinformation/unreliable news websites. While only 0.38% of articles on average per domain from our set of misinformation websites were classified as machine-generated in January 2022, 0.91% of articles on average from our set of mainstream/reliable websites were classified as machine-generated. This result is consistent with prior observations that many news websites have begun to use automated services to write quick, often financial-related articles (Section 2). For example, the beginning of one of the articles from Reuters (Figure 2) classified by our system as being machine-generated simply contained simple information about the direction of particular markets and funds.
@@
Figure 4: The number of articles across our set of 3,074 domains that contained a common ChatGPT error message over time.
However, despite reliable/mainstream websites initially having higher levels of synthetic text, misinformation websites had marked increases in levels of machine-generated content during 2022 and 2023 (Figure 5). While between January 1, 2022, and May 1, 2023, reliable/mainstream news websites had a 55.4% relative increase (0.51% absolute percentage increase) in their levels of synthetic content, misinformation websites had a 457% relative increase (1.75% absolute percentage increase). Starting from a lower base, we thus see a substantial increase in the number of synthetic articles from unreliable/misinformation websites.
Furthermore, as seen in Figure 3, we further observe that an increasing number of news outlets published at least one synthetic article within any given 30-day time frame. Across our period of study, the number of mainstream websites that published at least one synthetic article increased from 694 (35.4% of mainstream websites) in January 2022 to 934 (47.7%) in April 2023. Similarly, the number of misinformation websites that published at least one synthetic article increased from 119 (10.7% of misinformation websites) to 190 (17.1%).
@@
All $1.75\%$ $457\%$ $0.51\%$ $55.4\%$
Rank $<$ 10K $0.74\%$ $191\%$ $0.35\%$ $26.6\%$
10K $<$ Rank $<100$K $1.54\%$ $202\%$ $0.30\%$ $30.3\%$
100K $<$ Rank $<1$M $1.30\%$ $350\%$ $0.23\%$ $28.3\%$
1M $<$ Rank $<10$M $1.50\%$ $588\%$ $0.12\%$ $13.4\%$
Rank $>10$M $3.27\%$ $728\%$ $2.27\%$ $376\%$
@@
To confirm these initial findings, we further examine the increase in common idiosyncratic error messages often returned by ChatGPT. Specifically using a list of error messages including “my cutoff date in September 2021”, “as an AI language model”, and “I cannot complete this prompt” that the company News Guard (Sadeghi and Arvanitis 2023) has used to detect AI-generated websites, we gather every article among our 15.90 million articles that utilized such message: altogether 570 articles from 280 domains. Amongst these websites, the top domains of these articles included forbes.com (32 articles), dailymail.co.uk (29), fairobserver.com (19), theregister.com (13), and patheos.com (13).
@@
opensecrets.org 42.5% $<$100K thelist.com 26.0% $<$100K
theodysseyonline.com 26.2% $<$1M bjreview.com 26.0% $<$10M
logically.ai 16.8% $<$10M thefrisky.com 23.6% $<$1M
egypttoday.com 15.3% $<$1M northkoreatimes.com 23.1% $>$10M+
sourcewatch.org 13.9% $<$1M egypttoday.com 20.9% $<$1M
thequint.com 9.85% $<$100K waynedupree.com 20.1% $<$1M
bleacherreport.com 9.81% $<$10K ancient-origins.net 15.3% $<$100K
africanews.com 9.31% $<$1M entrepreneur.com 14.9% $<$100K
worldpopulationreview.com 9.27% $<$100K logically.ai 14.6% $<$10M
laweekly.com 9.14% $<$1M bignewsnetwork.com 14.5% $<$1M
Table 6: Websites with the largest percentage of synthetic content in January 2022 and in April 2023.
@@
We finally note that we observe a small but noticeable dip in the percentage and amount of synthetic content in Figures 5 (particularly among misinformation websites) and 4 between February and March 2023. We find that for Figure 5, unreliable websites such as pluralist.com, helsinkitimes.fi, and thelist.com, in particular, drove the initial increase in machine-generated content in January and February 2022, before dramatically decreasing their amount of synthetic content in the following month. Examining Google trends data, we also observe that ChatGPT experienced a noticeable dip/decline (from 87% of peak search traffic on February 5 to 75% peak search traffic on March 5) in popularity in the United States during this period perhaps (but not definitively) explaining this small decline.
@@
Examining these increases within particular brackets of popularity, we see (as pictured in Figure 5 and calculated in Table 5) that the least popular websites saw the largest percentage increase in the use of synthetic articles. For both unreliable/misinformation and reliable/mainstream categories, we observe that for websites that rank $>$10M+ in popularity, the percentage of their articles that were synthetic increased by 3.27% (728% relatively) and 2.27% (376%) on average, respectively. By contrast, among the most popular misinformation/unreliable websites (e.g., breitbart.com, zerohedge.com) and mainstream/reliable websites (e.g., cnn.com, foxnews.com), synthetic articles enjoyed only a smaller 0.74% (191% relatively) and a 0.35% (26.6%) increase overall. Indeed, calculating the websites with the most machine-generated content, we again observe in Table 6 that the websites that had the largest amounts of synthetic content were all fairly small or unpopular small.
@@
To identify the topics within our identified set of machine-generated articles, we train a DeBERTa-based classifier to identify the topic of an article based on its text. As training data, we utilize the News Catcher Topic Labelled dataset, [^13] https://www.kaggle.com/datasets/kotartemiy/topic-labeled-news-dataset which contains topic labels for 106,395 different articles as belonging to 8 different categories {Business, Entertainment, Health, US/Nation, Science, Sports, Technology, and World}. We note while the original dataset only contained the title of each article, the dataset also included the original URL. As such, using the method outlined in section 4, we gather the set of articles listed in the dataset and subsequently train a DeBERTa-based classifier to correctly label articles based on their content. We note that a significant portion of these URLs were not available; as a result, we trained our model on a subset of 79,000 articles from the original dataset, further removing articles that were less than 1000 characters. Keeping out a 10% of this dataset as a test dataset, upon training, we achieve a 0.819 $F_{1}$ score an average of 0.819 precision across the eight categories. Once trained, we finally categorize the topic of each of the 15.90 million articles within our dataset.
@@
Entertainment 0.71 Science 1.64
Business 0.63 Sports 0.40
Health 2.11 Technology 0.25
Nation 0.77 World 1.47
@@
Plotting the proportion of each topic amongst synthetic articles from misinformation websites, as seen in Figure 5(a), a significant portion of synthetic articles from misinformation websites concerned World affairs, Nation/US-current events, Science, and Entertainment. For example, among our set of synthetic articles from misinformation websites, we identify a variety of articles about concerns about tensions between Russia and Ukraine, COVID-19 vaccines, and updates about the love life of Ed Sheeran. Calculating the odds ratio between the number of synthetic and human-written articles for each of our topic categories, as seen in Table 7, among our selection of misinformation websites, relative to their own topic proportions, misinformation websites were most likely to utilize synthetic articles for Health and Science related topics. This suggests that misinformation websites have proportionally utilized synthetic articles for both mundane topics like Entertainment and and more serious topics such as health (Peiser 2019).
@@
Entertainment 0.60 Science 1.47
Business 1.51 Sports 0.81
Health 0.98 Technology 0.66
Nation 1.15 World 0.81
@@
Finally, calculating the odds ratio (Table 9) between the rates of usage of synthetic articles per category, we further observe that misinformation websites, throughout our period of study were more likely to utilize synthetic articles on topics related to Entertainment, Entertainment, and Science, and World affairs. In contrast, mainstream websites were more likely to utilize synthetic articles for Business, Technology (very small proportion), and Sports. We observe similar proportions of US/Nation topics between misinformation and mainstream websites.
@@
Entertainment 2.84 Science 1.85
Business 0.13 Sports 0.14
Health 1.40 Technology 0.13
Nation 1.01 World 2.80

### 20 Mar 2024 v5

@@ Abstract
As large language models (LLMs) like ChatGPT have gained traction, an increasing number of news websites have begun utilizing them to generate articles. However, not only can these language models produce factually inaccurate articles on reputable websites but disreputable news sites can utilize LLMs to mass produce misinformation. To begin to understand this phenomenon, we present one of the first large-scale studies of the prevalence of synthetic articles within online news media. To do this, we train a DeBERTa-based synthetic news detector and classify over 15.46 million articles from 3,074 misinformation and mainstream news websites. We find that between January 1, 2022, and May 1, 2023, the relative number of synthetic news articles increased by 57.3% on mainstream websites while increasing by 474% on misinformation sites. We find that this increase is largely driven by smaller less popular websites. Analyzing the impact of the release of ChatGPT using an interrupted-time-series, we show that while its release resulted in a marked increase in synthetic articles on small sites as well as misinformation news websites, there was not a corresponding increase on large mainstream news websites.
@@ 1 Introduction
In this work, we present a large-scale study of the relative increase in machine-generated/synthetic articles from 3,074 news websites (1,059 misinformation/unreliable websites and 2,015 mainstream/reliable news websites) between January 1, 2022, and May 1, 2023. To do this, we utilize data from 19 models, as well as adversarial data from article perturbation/re-writes and paraphrases, to train a DeBERTa-based model (He et al. 2021) to detect English-language synthetic news articles. We subsequently benchmark this classifier on eight test sets of machine-generated news articles, including two from real-world companies (Pu et al. 2022) and one from an independent collection of human-written real-world articles. Across these test datasets, our model, at a false positive rate (FPR) of 1%, achieves an average precision score of 0.992. With this trained model, we classify over 15.46M articles published between January 1, 2022, and May 1, 2023, from our set of 3,074 news websites. [^1] We release the weights of our model and the URLs used in this study at https://github.com/hanshanley/machine-made-media.
We find that among reliable/mainstream news websites, synthetic articles increased in prevalence by 57.3% (0.88% of news articles in January 2022 to 1.39% in May 2023) while among unreliable/misinformation websites, the prevalence increased by 474% (0.39% of news articles in January 2022 to 2.22% in May 2023). Examining the content of synthetic articles, we find that while mainstream/reliable news websites have largely utilized synthetic articles to report on financial and business-related news, misinformation/unreliable news websites have reported on a wide range of topics ranging from world affairs (e.g., the Russo-Ukrainian War) to human health (e.g., COVID-19). Examining the impact of ChatGPT on the prevalence of synthetic content, we further find that its release coincided with significant increases in machine-generated articles on misinformation websites and unpopular mainstream news websites.
@@ 3 Detecting Machine-Generated Articles
We finally note that, as found in prior work (Pu et al. 2022; Uchendu et al. 2021; Zellers et al. 2019), machine-generated news articles are often shorter in length than human-written articles. While training, to ensure that our models do not simply distinguish between longer human-written articles and those generated by generative transformers by their different lengths, we ensure that our machine-generated and human-written articles are of similar lengths (median training synthetic article length of 210 words and median training human article length of 224 words). Furthermore, as found by past work, predictions for texts, particularly short texts, tend to be unreliable (Kirchner et al. 2023; Zellers et al. 2019; Pu et al. 2022); conversely, as shown by Sadasivan et al. (2023), as the lengths of texts increase the variance between human and machine-generated texts increases. As such, for our training and our generated test data (GPT 3.5 dataset), we exclude texts shorter than 1,000 characters (140 words) (OpenAI 2022). We note as a result, we do not use every trained model’s articles from the Turing Benchmark; given that WMT-20/FAIR articles within this dataset are all shorter than 1000 characters, we do not include them within our training dataset. Altogether our training dataset thus includes data from 19 different models (18 from Turing Benchmark and GPT-3.5 davinici).
@@
Constructing Paraphrased Synthetic Articles. To paraphrase each of the machine-generated articles within our dataset, we use the approach outlined by Krishna et al. (2023). Specifically, as in their work, we utilize Dipper, a version of the T5 generative model fine-tuned on paragraph-level paraphrases, that outputs paraphrased versions of the inputted text. We use the default and recommended parameters [^5] We use a lexical diversity parameter of 60. For more details on the Dipper model see Krishna et al. (2023) as in Krishna et al. to paraphrase a portion of the text within our original training dataset as well as our GPT-3.5 test dataset (Krishna et al. 2023). We note that after paraphrasing our datasets, we again filter to ensure all articles utilized for training contain at least 1000 characters (Table 2). We annotate training and test datasets containing articles paraphrased with Dipper with the suffix Para. After paraphrasing we still consider these articles to be synthetic.
@@ 4 News Dataset and Classification Pipeline
Website List. Between January 1, 2022, and May 1, 2023, we gather all articles published from 3,074 news websites. [^10] We note that while this study focuses on the release of ChatGPT as a possible focal point, our data collection for this project actually began in January 2022. Our list of websites consists of domains labeled as “news” by Media Bias Fact Check [^11] https://mediabiasfactcheck.com/ and by prior work (Hanley, Kumar, and Durumeric 2023). Within our list of news sites, we differentiate between “unreliable news websites” and “reliable news websites.” Our list of unreliable news websites includes 1,059 domains labeled as “conspiracy/pseudoscience” by mediabiasfactcheck.com as well as those labeled as “unreliable news”, misinformation, or disinformation by prior work (Hanley, Kumar, and Durumeric 2023; Barret Golding 2022; Szpakowski 2020). Our set of “unreliable” or misinformation news websites includes websites like realjewnews.com, davidduke.com, thegatewaypundit.com, and breitbart.com. We note that despite being labeled unreliable every article from each of these websites is not necessarily misinformation.
Our set of “reliable”/mainstream news websites consists of the news websites that were labeled as belonging to the “center”, “center-left”, or “center-right” by Media Bias Fact Check as well as websites labeled as “reliable” or “mainstream” by other works (Hanley, Kumar, and Durumeric 2023; Barret Golding 2022; Szpakowski 2020). This set of “reliable news websites” includes websites like washingtonpost.com, reuters.com, apnews.com, cnn.com, and foxnews.com. Altogether after removing duplicates and unavailable websites, we scraped 2,015 “reliable news” or mainstream websites.
@@
Given that many of our websites (e.g., cnn.com) have multilingual options, we use the Python langdetect library to filter out all non-English articles. To prepare data for classification, we remove boilerplate language using the Pythonjustext library and then remove URLs, emojis, and HTML tags. Further, to ensure the reliability of our classifications, we only classify news articles that are at least 1000 characters (approximately 140 words) long. Altogether, from our selection of 3,074 websites, we gathered 15.46M articles (12.06M from mainstream websites and 3.39M from misinformation websites) that were published between January 1, 2022, and May 1, 2023. Finally, we utilize our DeBERTa+Pert+Para model at a softmax classification threshold of 0.98 to classify each article as either human-written or machine-generated. Classifying all 15.46M articles took approximately 65.8 hours using an Nvidia RTX A6000 GPU.
@@
With the rise of LLMs, many companies have widely scraped and gathered data from websites to fuel their models (Schappert 2023). As a result, websites ranging from Twitter to Reddit have begun to set up restrictions to ensure the privacy of their users and to protect their content from being used in other private companies’ generative models. While we do not train a generative model that could artificially produce convincing and seemingly unique reproductions of the texts that we utilize, we note the concern that our work raises. Our work, however, only studies the texts of our set of 15.46 million articles and classifies them as machine-generated or human-written. We do not seek to generate summaries or artificial rewrites of this content. In terms of web crawling for this data, as noted elsewhere (Singrodia, Mitra, and Paul 2019; Hanley, Kumar, and Durumeric 2023; Smith et al. 2013), website crawling and scraping remain pivotal for understanding and documenting what occurs on the Internet. Without scraping, understanding trends and how the Internet could potentially affect real life becomes impossible. As decided in Van Burn v. United States, publically accessible information can be legally scraped as long as it is done ethically and does no harm to the site (Emily R. Lowe and Katrina Slack 2022). As such, we collect only publicly available data from our set of websites and follow the best practices for web crawling as in Acar et al. (2014). We limit the load that each news site experiences by checking for new articles daily at a maximum rate of one request every 10 seconds. The hosts that we scan from are identifiable through WHOIS, reverse DNS, and an HTTP landing page explaining how to reach us if they would like to be removed from the study. During our crawling period, we received no requests from websites to opt out.
@@ 5 The Rise of Machine-Generated Media
Large-Scale Trends in Machine-Generated Media. To begin, we plot the average percentage of synthetic news articles per website across our dataset between January 1, 2022, and May 1, 2023, in Figure 1. In aggregate, across all 3,074 sites, we see that 1.07% of all articles published in January 2022 (12,984 of 1,213,983 articles) were synthetically generated. However, by May 2023, the fraction of synthetic articles nearly went up to 1.78% (25,561 of 1,439,812 articles), a 66.0% relative increase (nearly doubling in raw amount).
We observe that our set of reliable/mainstream websites typically had a greater percentage of synthetic articles at the beginning of 2022 compared with misinformation/unreliable news websites. While only 0.39% of articles on average per domain from our set of misinformation websites were classified as machine-generated in January 2022, 0.88% of articles on average from our set of mainstream/reliable websites were classified as machine-generated. This result is consistent with prior observations that many news websites have begun to use automated services to write quick, often financial-related articles (Section 2). For example, the beginning of one of the articles from Reuters (Figure 2) classified by our system as being machine-generated simply contained simple information about the direction of particular markets and funds.
@@
Figure 4: The number of articles that contained a common ChatGPT error message over time.
However, despite reliable/mainstream websites initially having higher levels of synthetic text, misinformation websites had marked increases in levels of machine-generated content during 2022 and 2023 (Figure 5). While between January 1, 2022, and May 1, 2023, reliable/mainstream news websites had a 57.3% relative increase (0.51% absolute percentage increase) in their levels of synthetic content, misinformation websites had a 474% relative increase (1.85% absolute percentage increase). Starting from a lower base, we thus see a substantial increase in the number of synthetic articles from unreliable/misinformation websites.
Furthermore, as seen in Figure 3, we further observe that an increasing number of news outlets published at least one synthetic article within any given 30-day time frame. Across our period of study, the number of mainstream websites that published at least one synthetic article increased from 697 (34.6% of mainstream websites) in January 2022 to 940 (46.6%) in April 2023. Similarly, the number of misinformation websites that published at least one synthetic article increased from 110 (10.4% of misinformation websites) to 179 (16.9%).
@@
All $1.85\%$ $474\%$ $0.51\%$ $57.3\%$
Rank $<$ 10K $0.70\%$ $175\%$ $0.36\%$ $27.0\%$
10K $<$ Rank $<100$K $1.77\%$ $221\%$ $0.26\%$ $27.3\%$
100K $<$ Rank $<1$M $1.38\%$ $349\%$ $0.23\%$ $30.1\%$
1M $<$ Rank $<10$M $1.55\%$ $646\%$ $0.14\%$ $15.4\%$
Rank $>10$M $3.42\%$ $736\%$ $2.13\%$ $423\%$
@@
To confirm these initial findings, we further examine the increase in common idiosyncratic error messages often returned by ChatGPT. Specifically using a list of error messages including “my cutoff date in September 2021”, “as an AI language model”, and “I cannot complete this prompt” that the company News Guard (Sadeghi and Arvanitis 2023) has used to detect AI-generated websites, we gather every article among our 15.46 million articles that utilized such message: altogether 570 articles from 280 domains. Amongst these websites, the top domains of these articles included forbes.com (32 articles), dailymail.co.uk (29), fairobserver.com (19), theregister.com (13), and patheos.com (13).
@@
opensecrets.org 42.5% $<$100K china.org.cn 34.9% $<$1M
theodysseyonline.com 26.2% $<$1M globaltimes.cn 26.3% $<$100K
logically.ai 17.2% $<$10M thelist.comm 26.0% $<$1M
china.org.cn 16.3% $<$1M bjreview.com 26.0% $>$10M+
globaltimes.cn 16.0% $<$100K thefrisky.com 23.6% $<$1M
egypttoday.com 15.0% $<$1M northkoreatimes.com 23.1% $>$10M+
sourcewatch.org 14.4% $<$1M egypttoday.com 21.0% $<$1M
bleacherreport.com 9.84% $<$100K waynedupree.com 20.1% $<$1M
thequint.com 9.81% $<$10K ancient-origins.net 15.3% $<$100K
africanews.com 9.64% $<$1M entrepreneur.com 15.0% $<$100K
Table 6: Websites with the largest percentage of synthetic content (with at least 100 articles in that month) in January 2022 and in April 2023.
@@
We finally note that we observe a small but noticeable dip in the percentage and amount of synthetic content in Figures 5 (particularly among misinformation websites) and 4 between February and March 2023. We find, as seen in Figure 5, that unreliable websites such as foreignpolicyi.org, prophecynewswatch.com, and awarenessact.com, in particular, drove the initial increase in machine-generated content in January and February 2022, before dramatically decreasing their amount of synthetic content in the following month. Examining Google trends data, we also observe that ChatGPT experienced a noticeable dip/decline (from 87% of peak search traffic on February 5 to 75% peak search traffic on March 5) in popularity in the United States during this period perhaps (but not definitively) explaining this small decline.
@@
Examining these increases within particular brackets of popularity, we see (as pictured in Figure 5 and calculated in Table 5) that the least popular websites saw the largest percentage increase in the use of synthetic articles. For both unreliable/misinformation and reliable/mainstream categories, we observe that for websites that rank $>$10M+ in popularity, the percentage of their articles that were synthetic increased by 3.42% (736% relative increase) and 2.13% (423%) on average, respectively. By contrast, among the most popular misinformation/unreliable websites (e.g., breitbart.com, zerohedge.com) and mainstream/reliable websites (e.g., cnn.com, foxnews.com), synthetic articles had a smaller 0.70% (175% relative increase) and a 0.36% (27.0%) increase overall. Indeed, calculating the websites with the most machine-generated content, we again observe in Table 6 that the websites that had the largest amounts of synthetic content were all fairly small or unpopular small.
@@
To identify the topics within our identified set of machine-generated articles, we train a DeBERTa-based classifier to identify the topic of an article based on its text. As training data, we utilize the News Catcher Topic Labelled dataset, [^13] https://www.kaggle.com/datasets/kotartemiy/topic-labeled-news-dataset which contains topic labels for 106,395 different articles as belonging to 8 different categories {Business, Entertainment, Health, US/Nation, Science, Sports, Technology, and World}. We note while the original dataset only contained the title of each article, the dataset also included the original URL. As such, using the method outlined in section 4, we gather the set of articles listed in the dataset and subsequently train a DeBERTa-based classifier to correctly label articles based on their content. We note that a significant portion of these URLs were not available; as a result, we trained our model on a subset of 79,000 articles from the original dataset, further removing articles that were less than 1000 characters. Keeping out a 10% of this dataset as a test dataset, upon training, we achieve a 0.819 $F_{1}$ score an average of 0.819 precision across the eight categories. Once trained, we finally categorize the topic of each of the 15.46 million articles within our dataset.
@@
Entertainment 0.68 Science 1.58
Business 0.61 Sports 0.23
Health 2.06 Technology 0.21
Nation 0.77 World 1.56
@@
Plotting the proportion of each topic amongst synthetic articles from misinformation websites, as seen in Figure 5(a), a significant portion of synthetic articles from misinformation websites concerned World affairs, Nation/US-current events, Science, and Entertainment. For example, among our set of synthetic articles from misinformation websites, we identify a variety of articles about concerns about tensions between Russia and Ukraine, COVID-19 vaccines, and updates about the love life of Ed Sheeran. Calculating the odds ratio between the number of synthetic and human-written articles for each of our topic categories, as seen in Table 7, among our selection of misinformation websites, relative to their own topic proportions, misinformation websites were most likely to utilize synthetic articles for Health and Science related topics. This suggests that misinformation websites have proportionally utilized synthetic articles for both mundane topics like Entertainment and more serious topics such as Health (Peiser 2019).
@@
Entertainment 0.59 Science 1.48
Business 1.53 Sports 0.85
Health 0.89 Technology 0.66
Nation 1.14 World 0.80
@@
Finally, calculating the odds ratio (Table 9) between the rates of usage of synthetic articles per category between mainstream and misinformation websites, we further observe that misinformation news and mainstream websites, throughout our period of study were more likely to utilize synthetic articles on topics related to Entertainment, Health, and Science, and World affairs. In contrast, mainstream websites were more likely to utilize synthetic articles for Business, Technology (very small proportion), and Sports. We observe similar proportions of US/Nation topics between misinformation and mainstream websites.
@@
Entertainment 2.96 Science 1.91
Business 0.13 Sports 0.06
Health 1.67 Technology 0.11
Nation 1.02 World 2.75